OpenStudy (anonymous):

Solve. (x + 2)2 = 1 help?

OpenStudy (anonymous):

it become 2x + 4 = 1 then you subtract both sides 2x + 4 - 4 = 1 - 4 and 2x = -3 x = 1.5

OpenStudy (anonymous):

these are the answer choices: x = -3 and x = -1 x = 3 and x = 1 x = -1 and x = 4 x = -4 and x = -3

OpenStudy (anonymous):

wait is the question (x + 2)^2 = 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats why

OpenStudy (anonymous):

haha yeah can you still help?

OpenStudy (anonymous):

yup its the first one

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

do you want me to explain?

OpenStudy (anonymous):

yes if you can that'd be great!

OpenStudy (anonymous):

ok you first have to make the equation into (x + 2)(x + 2) = 1 then you will get x^2 +2x + 2x + 4 = 1 then you will get x^2 + 4x + 4 = 1 so you can just subtract the 1 into the equation x^2 + 4x + 3 = 0 then the equation turns into (x + 3)(x + 1) so the x = -3 and -1

OpenStudy (anonymous):

THANKS!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yw